OpenStudy (itdontmatter):

Simplify the expression below. 3 · -8 + 5 - 4 ÷ 2

OpenStudy (kenrick143):

Familiar with PEMDAS?

Arcadiouse (arcadiouse):

Okay, do division and multiplication first. So now we have -24 + 5 -2

OpenStudy (itdontmatter):

yes

Arcadiouse (arcadiouse):

-24 + 5 = -19 -19 - 2 = -21

Arcadiouse (arcadiouse):

So your equation is simplified to -21

OpenStudy (kenrick143):

Where did you get -29?

OpenStudy (kenrick143):

I mean -19

OpenStudy (itdontmatter):

thank you

Arcadiouse (arcadiouse):

Because -24 + 5 = -19

Arcadiouse (arcadiouse):

The rule of negatives...

OpenStudy (kenrick143):

-4÷2 = -2

Arcadiouse (arcadiouse):

Not a negative...

OpenStudy (kenrick143):

Ahh.. I misunderstood apologies.

Arcadiouse (arcadiouse):

No problem.

OpenStudy (kenrick143):

Not a negative? @Arcadiouse

Arcadiouse (arcadiouse):

Yeah, the 4 is not a negative.

OpenStudy (kenrick143):

Its a negative since its minus 4

OpenStudy (kenrick143):

The answer is -2 where its 5-2